OpenStudy (anonymous):

use substitution to determine whether the point is on the line. x+6y=40 ; A (-10,5)

OpenStudy (anonymous):

Just plug the points into your equation. A pair of coordinates is always given (x,y), so -10 will be substituted for x and 5 will be substituted for y in the equation. So -10 + 6*5 = 40. If this is true, the point is on the line. If it's not true, the point is not on the line. You should get 20=40, which is not true, so in this problem the point is not on the line.

OpenStudy (anonymous):

ok. Thankyou so much. I understand now. WOW. your better than my math teacher.

OpenStudy (anonymous):

\[x+6y=40\] at point A (10,5). Remember that the point A (10,5) means that A's x value = 10, and A's y value = 5. So, replace the x & y in your equation with 10 & 5. If the left side of your equation equals the right side the point A is on the line \[x+6y=40\].

OpenStudy (anonymous):

Thanks guys! Really helped.